# 3DGS-DET: Empower 3D Gaussian Splatting with Boundary Guidance and Box-Focused Sampling for 3D Object Detection

### v1

@@ 3.2 Proposed Basic Pipeline of 3DGS for 3D Object Detection
In this section, we build our basic pipeline by directly utilizing the original 3D Gaussian Splatting (3DGS) for 3D Object Detection (3DOD) without any further improvement. As depicted in the top row of Fig. 2, we train the 3DGS representation of the input scene using posed images, denoted as $G=\left\{(\bm{\mu}_{i},\bm{S}_{i},\bm{R}_{i},\bm{c}_{i},\bm{\alpha}_{i})\right\}_{i=1}^{N}$. Given that the number of Gaussian blobs $N$ is too large for them to be input into the detector, we perform random sampling to select a subset of Gaussian blobs, denoted as $\hat{G}=\left\{(\bm{\mu}_{i},\bm{S}_{i},\bm{R}_{i},\bm{c}_{i},\bm{\alpha}_{i})\right\}_{i=1}^{M}$, where $M<N$. We then concatenate the attributes of the Gaussian blobs along the channel dimension as follows:
@@
$\hat{G}_{\text{input}}=\text{Concat}(\bm{\mu}_{i},\bm{S}_{i},\bm{R}_{i},\bm{c}_{i},\bm{\alpha}_{i})\quad\forall i\in\{1,\ldots,M\}.$ (5)
@@
This concatenated representation $\hat{G}_{\text{input}}$ is then fed into the subsequent detection tool. Note that since 3DGS is an explicit 3D representation, $\hat{G}_{\text{input}}$ can be utilized with any point-cloud-based detector by retraining the detector model on 3DGS representation. In our study, the research focus is on enhancing 3DGS for 3DOD in general, rather than designing a specific detector. Therefore, we utilize the existing work (Rukhovich et al., 2022a) as the detection tool. The final detection predictions are obtained as follows:
$P=\mathrm{F}(\hat{G}_{\text{input}})=\left(\bm{p},\bm{z},\bm{b}\right),$ (6)
@@
where $\mathrm{F}$ denotes the detector tool and $P$ represents the predictions, including classification probabilities $\bm{p}$, centerness $\bm{z}$, and bounding box regression parameters $\bm{b}$.
The training loss (Rukhovich et al., 2022a) is defined as:
$\begin{split}L_{\text{det}}=\frac{1}{N_{\text{pos}}}\sum_{\hat{x},\hat{y},\hat{z}}\left(\mathds{1}_{\{p(\hat{x},\hat{y},\hat{z})\neq 0\}}L_{\text{reg}}(\hat{\bm{b}},\bm{b})+\mathds{1}_{\{p(\hat{x},\hat{y},\hat{z})\neq 0\}}L_{\text{cntr}}(\hat{\bm{z}},\bm{z})+L_{\text{cls}}(\hat{\bm{p}},\bm{p})\right),\end{split}$ (7)
@@
where the number of matched positions $N_{\text{pos}}$ is given by $\sum_{\hat{x},\hat{y},\hat{z}}\mathds{1}_{\{p(\hat{x},\hat{y},\hat{z})\neq 0\}}$. Ground truth labels are indicated with a hat symbol. The regression loss $L_{\text{reg}}$ is based on Intersection over Union (IoU), the centerness loss $L_{\text{cntr}}$ uses binary cross-entropy, and the classification loss $L_{\text{cls}}$ employs focal loss. Further details on the detection tool can be found in Rukhovich et al. (2022a).
Building upon this basic pipeline, we develop our method, 3DGS-DET, by introducing two novel designs to improve the 3DGS representation, as illustrated in the bottom row of Fig. 2. These designs are detailed in the following sections  Sec. 3.3 and Sec. 3.4.
@@ 3.3 Boundary Guidance
Given the fact that 3DGS reconstruction is derived from 2D images, we design the novel Boundary Guidance strategy by incorporating 2D Boundary Guidance to achieve a more suitable 3D spatial distribution of Gaussian blobs for detection. In this section, we present our Boundary Guidance strategy in detail. As illustrated in the top row of Fig. 3, to provide the guidance priors for 3DGS reconstruction, we first generate category-specific boundaries for posed images:
@@
where $\mathrm{H}_{\text{bd}}$ is the boundary generator, and $b_{\text{bd}}^{c}$ represents the binary boundary map for category $c$. If $b_{\text{bd}}^{c}(x,y)=1$, the pixel at $(x,y)$ belongs to the boundary for objects of category $c$. The set $C$ includes all categories. In practice, the operations of $\mathrm{H}_{\text{bd}}$ are as follows: we use Grounded SAM (Ren et al., 2024) to generate category-specific masks. Then, the Suzuki-Abe algorithm (Suzuki et al., 1985) is employed to extract the boundaries of these masks, along with category information. The category-specific boundaries are then overlaid on the posed images in different colors:
$I_{\text{bd}}(x,y)=I(x,y)\cdot\left(1-\sum_{c\in C}b_{\text{bd}}^{c}(x,y)\right)+\sum_{c\in C}b_{\text{bd}}^{c}(x,y)\cdot\text{color}(c),$ (9)
@@
To effectively reduce $L_{\text{render}}$ during training, it is crucial to ensure the rendering quality of boundaries and the multi-view stability of boundaries. In this way, the Boundary Guidance lead 3DGS to incorporate boundary prior information into the 3D space. As shown in Fig. 2 (better viewed when zoomed in), 3DGS trained with Boundary Guidance demonstrates improved spatial distribution of Gaussian blobs compared to those trained without it, without introducing additional learnable parameters.
@@ 3.4 Box-Focused Sampling
Considering that 2D images often include numerous background pixels, leading to densely reconstructed 3DGS with many noisy Gaussian blobs representing the background, negatively affecting detection. To reduce the excessive background blobs, in this section, we propose the Box-Focused Sampling strategy in detail. As depicted in the bottom row of Fig. 3, to provide priors for the following sampling, we utilize a 2D object detector to identify object bounding boxes:
@@
where $\mathrm{H}_{\text{bb}}$ is the box detector, and we select Grounding DINO (Liu et al., 2023) as the detector in our experiments. Here, $b_{\text{bb}}$ denotes the bounding box positions, and $p^{C}$ is the probability vector for the box belonging to each category in $C$. We define $p_{\max}=\max_{c\in C}p^{c}$ as the highest category probability for a given bounding box, which helps to establish object probability spaces in later step. Then, we project the 2D boxes into 3D space:
@@
where $F_{\text{ft}}$ is the projected 3D frustum from $b_{\text{bb}}$, and $K^{-1}$ is the inverse camera matrix used to map 2D bounding box corners $(x_{i},y_{i})$ and depth values $z_{\min}$ and $z_{\max}$ into 3D space. Next, we establish object probability spaces using $F_{\text{ft}}$ and $p_{\max}$. Specifically, for each bounding box, the maximum probability $p_{\max}$ models the likelihood of each Gaussian blob within the corresponding frustum being an object blob:
@@
where $p_{\text{obj}}(g_{i}\mid g_{i}\in F_{\text{ft}})$ indicates the probability of each Gaussian blob $g_{i}$ within frustum $F_{\text{ft}}$ being an object blob. To integrate priors from different view frustums, we select the maximum probability as the aggregated probability:
@@
where $p_{\text{agr}}(g_{i})$ is the aggregated probability for Gaussian blob $g_{i}$, and $V$ represents the set of all views. Gaussian blobs not belonging to any frustum are assigned a small probability $p_{\text{bg}}$, set to 0.01 in practice. In this way, we obtain the object probability spaces $P_{\text{obj}}$, where each Gaussian blob has an associated probability of being an object. We then perform independent probabilistic sampling based on $P_{\text{obj}}$ to achieve Box-Focused Sampling, resulting in the sampled Gaussian set $\hat{G}_{\text{bd}}^{\text{bs}}$ as:
@@
In this way, it allows object blobs to be better preserved due to their higher probabilities, while most background points, having lower probabilities, are effectively reduced. Then, based on $\hat{G}_{\text{bd}}^{\text{bs}}$, we proceed with the training of the detector, as formulated by Equ. 5-Equ. 7 as described in Sec. 3.2. As shown in Fig. 2, 3DGS sampled via Box-Focused Sampling retains more object blobs and reduces background noise.
@@ 4 Experiments
Table 1: Comparison of mAP@0.25 across different methods on ScanNet. The first block includes methods using non-view-synthesis representations, such as point cloud, RGB-D, and multi-view images. The second block includes methods utilizing view-synthesis representations (NeRF-based and our 3DGS-based method). Our 3DGS-DET significantly outperforms the NeRF-based method NeRF-Det by 6.6 points. For other representations, 3DGS-DET surpasses all methods except for the point-cloud-based methods, FCAF3D and CAGroup3D, which have inherent advantages by directly using sensor-captured 3D data, specifically point clouds, as input.
Methods cab bed chair sofa tabl door wind bkshf pic cntr
Seg-Cluster (Wang et al., 2018) 11.8 13.5 18.9 14.6 13.8 11.1 11.5 11.7 0.0 13.7
Mask R-CNN (He et al., 2017) 15.7 15.4 16.4 16.2 14.9 12.5 11.6 11.8 19.5 13.7
SGPN (Wang et al., 2018) 20.7 31.5 31.6 40.6 31.9 16.6 15.3 13.6 0.0 17.4
3D-SIS (Hou et al., 2019) 12.8 63.1 66.0 46.3 26.9 8.0 2.8 2.3 0.0 6.9
3D-SIS (w/ RGB) (Hou et al., 2019) 19.8 69.7 66.2 71.8 36.1 30.6 10.9 27.3 0.0 10.0
VoteNet (Qi et al., 2019) 36.3 87.9 88.7 89.6 58.8 47.3 38.1 44.6 7.8 56.1
FCAF3D (Rukhovich et al., 2022a) 57.2 87.0 95.0 92.3 70.3 61.1 60.2 64.5 29.9 64.3
CAGroup3D (Wang et al., 2022a) 60.4 93.0 95.3 92.3 69.9 67.9 63.6 67.3 40.7 77.0
ImGeoNet (Tu et al., 2023) 40.6 84.1 74.8 75.6 59.9 40.4 24.7 60.1 4.2 41.2
CN-RMA (Shen et al., 2024a) 42.3 80.0 79.4 83.1 55.2 44.0 30.6 53.6 8.8 65.0
ImVoxelNet (Rukhovich et al., 2022b) 30.9 84.0 77.5 73.3 56.7 35.1 18.6 47.5 0.0 44.4
NeRF-Det (Xu et al., 2023) 37.6 84.9 76.2 76.7 57.5 36.4 17.8 47.0 2.5 49.2
3DGS-DET (Our basic pipeline) 39.6 82.5 75.8 78.0 53.6 36.1 26.9 41.8 11.9 56.0
3DGS-DET (Our basic pipeline+BG) 38.9 83.5 81.7 82.6 54.4 36.2 26.0 39.6 13.5 52.8
3DGS-DET (Our basic pipeline+BG+BS) 44.1 82.7 81.7 79.6 56.0 35.4 27.6 45.2 17.3 61.9
Methods desk curt fridg showr toil sink bath ofurn mAP@0.25
Seg-Cluster (Wang et al., 2018) 12.2 12.4 11.2 18.0 19.5 18.9 16.4 12.2 13.4
Mask R-CNN (He et al., 2017) 14.4 14.7 21.6 18.5 25.0 24.5 24.5 16.9 17.1
SGPN (Wang et al., 2018) 14.1 22.2 0.0 0.0 72.9 52.4 0.0 18.6 22.2
3D-SIS (Hou et al., 2019) 33.3 2.5 10.4 12.2 74.5 22.9 58.7 7.1 25.4
3D-SIS (w/ RGB) (Hou et al., 2019) 46.9 14.1 53.8 36.0 87.6 43.0 84.3 16.2 40.2
VoteNet (Qi et al., 2019) 71.7 47.2 45.4 57.1 94.9 54.7 92.1 37.2 58.7
FCAF3D (Rukhovich et al., 2022a) 71.5 60.1 52.4 83.9 99.9 84.7 86.6 65.4 71.5
CAGroup3D (Wang et al., 2022a) 83.9 69.4 65.7 73.0 100.0 79.7 87.0 66.1 75.12
ImGeoNet (Tu et al., 2023) 70.9 33.7 54.4 47.5 95.2 57.5 81.5 36.1 54.6
CN-RMA (Shen et al., 2024a) 70.0 44.9 44.0 55.2 95.4 68.1 86.1 49.7 58.6
ImVoxelNet (Rukhovich et al., 2022b) 65.5 19.6 58.2 32.8 92.3 40.1 77.6 28.0 49.0
NeRF-Det (Xu et al., 2023) 52.0 29.2 68.2 49.3 97.1 57.6 83.6 35.9 53.3
3DGS-DET (Our basic pipeline) 69.8 36.7 38.3 55.3 93.5 64.0 80.8 37.5 54.3
3DGS-DET (Our basic pipeline+BG) 68.6 45.2 52.7 45.0 98.3 69.6 84.3 48.0 56.7
3DGS-DET (Our basic pipeline+BG+BS) 72.8 40.7 56.6 71.9 98.5 72.2 88.3 46.7 59.9 (+6.6)

### v2

@@ 3.2 Proposed Basic Pipeline of 3DGS for 3DOD
In this section, we build our basic pipeline by directly utilizing the original 3DGS for indoor 3D Object Detection (3DOD) without any further improvement. First, we train the 3DGS representation of the input scene using posed images, denoted as $G=\left\{(\boldsymbol{\mu}_{i},\boldsymbol{S}_{i},\boldsymbol{R}_{i},\boldsymbol{c}_{i},\boldsymbol{\alpha}_{i})\right\}_{i=1}^{N}$. Given that the number of Gaussian blobs $N$ is too large for them to be input into the detector, we perform random sampling to select a subset of Gaussian blobs, denoted as $\hat{G}=\left\{(\boldsymbol{\mu}_{i},\boldsymbol{S}_{i},\boldsymbol{R}_{i},\boldsymbol{c}_{i},\boldsymbol{\alpha}_{i})\right\}_{i=1}^{M}$, where $M<N$. We then concatenate the attributes of the Gaussian blobs along the channel dimension as follows:
@@
$\hat{G}_{\text{input}}=\text{Concat}(\boldsymbol{\mu}_{i},\boldsymbol{S}_{i},\boldsymbol{R}_{i},\boldsymbol{c}_{i},\boldsymbol{\alpha}_{i})\quad\forall i\in\{1,\ldots,M\}.$ (1)
@@
This concatenated representation $\hat{G}_{\text{input}}$ is then fed into the subsequent detection tool. Note that since 3DGS is an explicit 3D representation, $\hat{G}_{\text{input}}$ can be utilized with any point-cloud-based detector by retraining the detector model on 3DGS representation. In our study, the research focus is on enhancing 3DGS for 3DOD in general, rather than designing a specific detector. Therefore, we utilize the existing work [37] as the detection tool. The final detection predictions are obtained as follows:
$P=\mathrm{F}(\hat{G}_{\text{input}})=\left(\boldsymbol{p},\boldsymbol{z},\boldsymbol{b}\right),$ (2)
@@
where $\mathrm{F}$ denotes the detector tool and $P$ represents the predictions, including classification probabilities $\boldsymbol{p}$, centerness $\boldsymbol{z}$, and bounding box regression parameters $\boldsymbol{b}$. The training loss [37] is defined as:
$L_{\text{det}}=$ $\frac{1}{N_{\text{pos}}}\sum_{\hat{x},\hat{y},\hat{z}}\Big(\mathds{1}_{\{p(\hat{x},\hat{y},\hat{z})\neq 0\}}L_{\text{reg}}(\hat{\boldsymbol{b}},\boldsymbol{b})$
$+\mathds{1}_{\{p(\hat{x},\hat{y},\hat{z})\neq 0\}}L_{\text{cntr}}(\hat{\boldsymbol{z}},\boldsymbol{z})+L_{\text{cls}}(\hat{\boldsymbol{p}},\boldsymbol{p})\Big),$ (3)
@@
where the number of matched positions $N_{\text{pos}}$ is given by $\sum_{\hat{x},\hat{y},\hat{z}}\mathds{1}_{\{p(\hat{x},\hat{y},\hat{z})\neq 0\}}$. Ground truth labels are indicated with a hat symbol. The regression loss $L_{\text{reg}}$ is based on Intersection over Union (IoU), the centerness loss $L_{\text{cntr}}$ uses binary cross-entropy, and the classification loss $L_{\text{cls}}$ employs focal loss. Further details on the detection tool can be found in [37]. Building upon this basic pipeline, we develop our method, 3DGS-DET, by introducing two novel designs (Sec. 3.3 and Sec. 3.4) to improve the 3DGS representation, as illustrated in Fig. 2.
[…]
Figure 4: Qualitative Analysis of Box-Focused Sampling for 3DGS Blobs. Box-Focused Sampling significantly retains more object blobs and reduces noisy background blobs. Note that we visualize only the positions of the Gaussian blobs to highlight their spatial distribution, while omitting other attributes for clarity.
@@ 3.3 Boundary Guidance
Given the fact that 3DGS reconstruction is derived from 2D images, we design the novel Boundary Guidance strategy by incorporating 2D Boundary Guidance to achieve a more suitable 3D spatial distribution of Gaussian blobs for detection. In this section, we present our Boundary Guidance strategy. As illustrated in the blue block in the top row of Fig. 2, to provide the guidance priors for 3DGS reconstruction, we first generate category-specific boundaries for posed images:
@@
where $\mathrm{H}_{\text{bd}}$ is the boundary generator, and $b_{\text{bd}}^{c}$ represents the binary boundary map for category $c$. If $b_{\text{bd}}^{c}(x,y)=1$, the pixel at $(x,y)$ belongs to the boundary for objects of category $c$. The set $C$ includes all categories. In practice, the operations of $\mathrm{H}_{\text{bd}}$ are as follows: we use Grounded SAM [36] to generate category-specific masks. Then, the Suzuki-Abe algorithm [44] is employed to extract the boundaries of these masks, along with category information. The category-specific boundaries are then overlaid on the posed images in different colors:
$I_{\text{bd}}(x,y)=$ $I(x,y)\cdot\left(1-\sum_{c\in C}b_{\text{bd}}^{c}(x,y)\right)$
$+\sum_{c\in C}b_{\text{bd}}^{c}(x,y)\cdot\text{color}(c),$ (5)
@@
To effectively reduce $L_{\text{render}}$ during training, it is crucial to ensure the rendering quality of boundaries and the multi-view stability of boundaries. In this way, the Boundary Guidance lead 3DGS to incorporate boundary prior information into the 3D space. As shown in Fig. 3DGS-DET: Empower 3D Gaussian Splatting with Boundary Guidance and Box-Focused Sampling for Indoor 3D Object Detection, 3DGS trained with Boundary Guidance demonstrates improved spatial distribution of Gaussian blobs compared to those trained without it. Besides, Fig. 3 shows the rendered images from different views by 3DGS trained with Boundary Guidance. The category-specific boundaries are clearly rendered with preserved multi-view consistency, evidencing that the 3D representation has successfully embedded the priors from Boundary Guidance.
Methods cab bed chair sofa tabl door wind bkshf pic cntr desk curt fridg showr toil sink bath ofurn mAP@0.25
VoteNet 36.3 87.9 88.7 89.6 58.8 47.3 38.1 44.6 7.8 56.1 71.7 47.2 45.4 57.1 94.9 54.7 92.1 37.2 58.7
FCAF3D 57.2 87.0 95.0 92.3 70.3 61.1 60.2 64.5 29.9 64.3 71.5 60.1 52.4 83.9 99.9 84.7 86.6 65.4 71.5
CAGroup3D 60.4 93.0 95.3 92.3 69.9 67.9 63.6 67.3 40.7 77.0 83.9 69.4 65.7 73.0 100.0 79.7 87.0 66.1 75.12
ImGeoNet 40.6 84.1 74.8 75.6 59.9 40.4 24.7 60.1 4.2 41.2 70.9 33.7 54.4 47.5 95.2 57.5 81.5 36.1 54.6
CN-RMA 42.3 80.0 79.4 83.1 55.2 44.0 30.6 53.6 8.8 65.0 70.0 44.9 44.0 55.2 95.4 68.1 86.1 49.7 58.6
ImVoxelNet 30.9 84.0 77.5 73.3 56.7 35.1 18.6 47.5 0.0 44.4 65.5 19.6 58.2 32.8 92.3 40.1 77.6 28.0 49.0
NeRF-Det 37.6 84.9 76.2 76.7 57.5 36.4 17.8 47.0 2.5 49.2 52.0 29.2 68.2 49.3 97.1 57.6 83.6 35.9 53.3
NeRF-Det++ 36.1 82.9 74.9 79.1 57.0 37.3 24.9 54.6 2.4 51.7 72.2 25.5 58.7 51.5 92.7 50.8 82.2 35.1 53.9
3DGS-DET 44.1 82.7 81.7 79.6 56.0 35.4 27.6 45.2 17.3 61.9 72.8 40.7 56.6 71.9 98.5 72.2 88.3 46.7 59.9 (+6.0)
Table 1: Comparison of mAP@0.25 on the ScanNet dataset. The first two blocks include methods that utilize non-view-synthesis representations: the first block [33, 37, 46] features approaches based on point clouds, while the second block [45, 41, 38] focuses on multi-view image methods. The third block [52, 16] encompasses methods that employ view-synthesis representations, including NeRF-based approaches and our 3DGS-based method. Among non-view-synthesis representations, our approach clearly outperforms methods that utilize multi-view images (the second block), highlighting the superiority of our 3DGS reconstruction for detection with multi-view image inputs. In terms of view-synthesis representations, our 3DGS-DET surpasses the NeRF-based method NeRF-Det++ [16] by 6.0 points.
Methods cab fridg shlf stove bed sink wshr tolt bthtb oven dshw frplce stool chr tble TV sofa mAP@0.25
ImVoxelNet 32.2 34.3 4.2 0.0 64.7 20.5 15.8 68.9 80.4 9.9 4.1 10.2 0.4 5.2 11.6 3.1 35.6 23.6
NeRF-Det 36.1 40.7 4.9 0.0 69.3 24.4 17.3 75.1 84.6 14.0 7.4 10.9 0.2 4.0 14.2 5.3 44.0 26.7
NeRF-Det++ - - - - - - - - - - - - - - - - - 43.3
3DGS-DET (Ours) 45.2 84.4 33.3 41.4 87.3 75.5 67.6 87.2 90.8 74.3 6.0 56.4 26.3 70.3 60.6 0.7 81.8 58.2  (+14.9)
Table 2: Comparison of the ‘whole-scene’ performance [38, 52] on the ARKITScenes validation set. Our 3DGS-DET significantly outperforms NeRF-Det++ [16] by 14.9 points. Note that we follow the setup described in the NeRF-Det supplementary materials: ‘In our experiments, we utilize the subset of the dataset with low-resolution images’, considering it is the closest work to ours. Other methods that do not use the same setting are not listed in this table. NeRF-Det++ [16] does not report per-category performance in ARKITScenes. We report its overall mAP@0.25 following the original paper [16].
@@ 3.4 Box-Focused Sampling
Considering that 2D images often include numerous background pixels, leading to densely reconstructed 3DGS with many noisy Gaussian blobs representing the background, negatively affecting detection. To reduce the excessive background blobs, in this section, we propose the Box-Focused Sampling strategy in detail. As depicted in the orange block in the bottom row of Fig. 2, to provide priors for the following sampling, we utilize a 2D object detector to identify object bounding boxes:
@@
where $\mathrm{H}_{\text{bb}}$ is the box detector, and we select Grounding DINO [22] as the detector in our experiments. Here, $b_{\text{bb}}$ denotes the predicted bounding box positions, and $p^{C}$ is the predicted probability vector for the box belonging to each category in $C$. We define $p_{\max}=\max_{c\in C}p^{c}$ as the highest category probability for a given bounding box, which helps to establish object probability spaces in later step. Then, we project the 2D boxes into 3D space:
@@
where $F_{\text{ft}}$ is the projected 3D frustum from $b_{\text{bb}}$, and $K^{-1}$ is the inverse camera matrix used to map 2D bounding box corners $(x_{i},y_{i})$ and depth values $z_{\min}$ and $z_{\max}$ into 3D space. Next, we establish object probability spaces by $F_{\text{ft}}$ and $p_{\max}$. For each bounding box, the maximum probability $p_{\max}$ models the likelihood of Gaussian blobs within the corresponding frustum belonging to object blobs:
@@
where $p_{\text{obj}}(g_{i}\mid g_{i}\in F_{\text{ft}})$ denotes the probability of Gaussian blob $g_{i}$ within frustum $F_{\text{ft}}$ belonging to object blobs. To integrate priors from different view frustums, we select the maximum probability as the aggregated probability:
@@
where $p_{\text{agr}}(g_{i})$ is the aggregated probability for Gaussian blob $g_{i}$, and $V$ represents the set of all views. Gaussian blobs not belonging to any frustum are assigned a small probability $p_{\text{bg}}$, set to 0.01 in practice. In this way, we obtain the object probability spaces $P_{\text{obj}}$, where each Gaussian blob has an associated probability of belonging to an object. We then perform probabilistic sampling based on $P_{\text{obj}}$ to achieve Box-Focused Sampling, resulting in the sampled Gaussian set $\hat{G}_{\text{bd}}^{\text{bs}}$ as:
@@
In this way, it allows object blobs to be better preserved due to their higher probabilities, while most background points, having lower probabilities, are effectively reduced. Then, based on $\hat{G}_{\text{bd}}^{\text{bs}}$, we proceed with the training of the detector, as formulated by Equ. 1-Equ. 3 as described in Sec. 3.2. As shown in Fig. 4, 3DGS sampled via Box-Focused Sampling retains more object blobs and reduces background noise.